Project

General

Profile

Bug #1708

Muon Beamline Step IV prep list

Added by Nebrensky, Henry over 7 years ago. Updated over 7 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
13 July 2015
Due date:
08 September 2015
% Done:

100%

Estimated time:

Description

Issues that need sorting out before Step IV


Related issues

Precedes Operations - Bug #1788: Muon Beamline Step IV Prep listClosedNebrensky, Henry28 October 201516 February 2016

Actions
#1

Updated by Nebrensky, Henry over 7 years ago

Stuff that needs sorting out by 8th September 2015

Target

  • Light Loss (TGT) - stabilised in mild saturation, now falling again
  • Fibre connections beneath MLCR floor (TGT) - October shutdown
  • Fibre connections at Vac-Rack (TGT) - October shutdown
  • S3 and T2 - assembly of next spare (TGT) - pinch point is getting optics block off T1
  • Shifter "Target Drive" panel (CAM)
  • Target Alarm handler #1613 update 22 (CAM)
  • ISIS data items #1669 (CAM)
  • Telephone link #977
  • Activation Study - 18th October
  • Beambump tune done for cycle 2015/02

Conventional Magnets

  • Second chiller pump or hard limit in controls (VC175 )
    • need limit in control software (CAM)
  • Q9 Magnet-On lamps (DL)
  • Q4 oscillation below 10A and anomalous ramp rate (DL)
  • Q1 MPS start-up recalcitrance (DL)
  • Q1 alarm on voltage at zero amps (CAM)
  • Q4 Earth Leakage (MBC)
  • "All to 0A" controls issue
  • Water, temperatures and MPS water flows in ALH (CAM)
    • Q5 and Q6 temperature sensors swapped
  • POMPOM readout into EPICS (CAM)
    • POMPOMs in ALH?
  • Update magnet and POMPOM user guides (MBC)
  • Annual POMPOM radiation damage test (MBC)

Decay Solenoid

  • CV2 fix (DK) - replaced, testing
  • Overheating investigation (DK/DL)
  • 120A trip investigation (DK/DL)
  • Replace final battery cell (DL)
  • Diodes in BBU circuit (DL)
  • Linde PC install: awaiting validation and IOC (DK) - licence issues
    • MLCR monitor and KVM implications
  • Vac pump tables
  • ALH/archiver corrections #1710 (CAM)
  • Off-site EPICS Gateway/Archiver access #1460
  • Power control fix / "All to 0A" controls issue (CAM)
  • Expert-only mode
  • Magnet-On status lamp (DL) - needs synch access
  • Fix lamp for compressor bottle camera

Proton Absorber

  • Adjust Damping valves / Mechanical service: acquiring restrictor valves (GB)
  • Combine survey documents (MBC)

Beamstop

  • Replace Hose (MBC)
  • Frame release key - awaiting string
  • MICE-ISIS Safety mtg permission (MBC)
  • Translate Survey document (MBC)

General

  • BLOC training and experience (MBC)
    • Second beamline phone MBC
  • Shared services/environment (water, air, computing/networking, UPS ( #1213 ) in Alarm Handler and in Archiver (CAM)

Beyond the Muon Beamline

#2

Updated by Nebrensky, Henry over 7 years ago

  • Status changed from Open to Closed
  • % Done changed from 0 to 100

Close this one and create new issue #1788 for next year...

Also available in: Atom PDF